OpenStudy (anonymous):

Use the Substitution method to solve the system of equations. x + y = -4 x - y = 2

OpenStudy (anonymous):

Can you write a little more clearly? I think your equations are.. x+y=-4 x-y=2 well in order to use substitution we have to solve for a variable in one of the equations. Thus, isolating that variable. We can solve for x in equation 2 and get x=2+y We now plug this into the first equation for x 2+y+y=-4 2y=-6 y=-3 Now that you have y, plug it into the equation for x=2+y and solve for x

OpenStudy (anonymous):

how do you solve for x?

OpenStudy (anonymous):

well y=-3.. and we have an equation that says 2+y=x so we replace y=-3 and solve

OpenStudy (anonymous):

x+y = -4 x+(-3) = -4 x-3 = -4 x = -1

OpenStudy (anonymous):

thanks